OpenStudy (anonymous):

quadratic equation x(2x+4)=240

geerky42 (geerky42):

Factor 2 from (2x+4) out then divide both sides by 2. You should get x(x+2) = 120 x(x+2) - 120 = 0 x² + 2x - 120 = 0 Factor. (x+12)(x-10) = 0 x = 10, -12